OpenStudy (anonymous):

how is the derivative of 2x stated ?

OpenStudy (anonymous):

using the integral sign

OpenStudy (mertsj):



OpenStudy (anonymous):

what does the dx represent ?

OpenStudy (mertsj):

If y = 2x then \[\frac{dy}{dx}=2\] So \[dy=2dx\] And \[\int\limits dy=\int\limits 2x dx \]

OpenStudy (mertsj):

Whoops. Get rid of the x after the 2

OpenStudy (anonymous):

so it would just be 2dx

OpenStudy (kropot72):

The derivative of 2x is expressed as\[\frac{d(2x)}{dx}\]

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

thankyou both !